True or false?

(a) For an elementary step, the order of the reaction is equal to the molecularity of the reaction.

(b) For an elementary step, the order of the reaction is equal to the sum of the stoichiometric coefficients for the reaction.

(c) The order of an overall reaction is equal to the molecularity of the reaction.

(d) The molecularity of an overall reaction is equal to the sum of the stoichiometric coefficients for the reaction.

(e) The order of a reaction can be determined from the units of the rate constant for the reaction.

If you could explain why each is true or false it would be greatly appreciated!
Answer #1

statements A , B are true since elementary reaction means single step reaction. Order is equal to moleuclarity and is equal to sum of coeffcients of reactants when we consider slowest step in reaction mechanis. In elementary reaction we have only 1 step , hence order = molecularity = sum of coefficients of reactantas

C) False , since in overall reaction we consider slowest step and order is determinded from slowest step

D) True , moleuclarity depends on number of reactanats moleucles shown overall

E) True , we have relation rate constant units = ( M) ^1-n s-1

where is order , if 1st order n = 1 , then units of rate constant = s-1, if n = 2 then units = M-1s-1